OpenStudy (anonymous):

Jamal has a 60-month fixed installment loan, with a monthly payment of $310.50. The amount he borrowed was $11,500. Instead of making his 35th payment, Jamal is paying the remaining balance on the loan. How much interest will Jamal save?

OpenStudy (wolf1728):

I'm thinking you'd have to solve for the rate which is a hefty 20.905% Here's a calculator that will show interest payments by each month. http://www.1728.org/mortmnts.htm Granted I'm not telling you the exact process to calculate the answer but at least it will help to have an answer.

OpenStudy (wolf1728):

Okay well, here is the exact formula you will need: If he is going to make his 35th payment then that means he has already made 34 payments and the formula will be filled in like this: Balance= 11,500 * (1 + r)^34 -310.50 * ((1 + r)^34-1)/r where r = 20.905/1,200